A closed organ pipe can vibrate at a minimum frequency of 500 Hz. Find the length of the tube. Speed of sound in air = 340 m s-1.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. V = 340 m/s Distances between two nodes or antinodes ⇒ λ/4 = 25 cm ⇒ λ = 100 cm = 1 m ⇒ n = v/λ = 340 Hz.